• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kalyan

    Log entry time 20:47:12 on November 11, 2008

    Entry number 248234

    This entry is a followup to: 248104

    Followups:

    keyword=DAQ activity today

    1. I removed all the delay on T3 at the coincidence circuit and also reduced the T1 delay to make the T5. As before, all the triggers come within about 20ns at the trigger supervisor.

    2. The gates have been adjusted for the BigBite ADCs. TDC timing will change on all the BigBite signals due to this change. Wirechamber t0s will have to be re-calibrated with T1.

    3. Checked to make sure the re-timing is ok for the L-HRS.

    4. Shower channels 19.4 and 22.5 signals which were showing low counts in beam actually looks ok with cosmics(not any different from others).

    5. Alex suggested not to do re-timing at this point. One thing at a time...I think we will look at the data tonight and then decide later.



    A copy of this log entry has been emailed to: jiang,brads,camsonne,rom,vasulk